# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

### v3

@@ 8 Conclusion
We have presented Orthogonalized Policy Optimization (OPO), a framework grounded in two independent geometries: an $\alpha$-divergence escort projection that shapes sampling, and a Euclidean mirror map (Pearson $\chi^{2}$) that shapes optimization. The same loss emerges from two complementary derivations—mirror descent and ratio-space regression—confirming it is uniquely determined by the geometry choice. The resulting objective has a constant Hessian, linear non-saturating gradients, and global contraction, addressing a key limitation of KL-based methods in high-confidence reasoning tasks. We believe OPO provides a principled foundation for future work on robust and scalable LLM alignment.
@@ A.1 Log-Ratio Approximation Error
Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ be the log-ratio, and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$ be the exact ratio deviation. Let $\delta=\|\Delta_{\theta}\|_{\infty}$. For sufficiently small $\delta$ (in particular $\delta<1$, which is enforced by the $\chi^{2}$ trust-region penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ bounding $\|\Delta\|_{\infty}$):

### v4

@@ 11 Conclusion
We have presented Orthogonalized Policy Optimization (OPO), grounded in a unified work–dissipation principle: the policy update maximizes external work from an $\alpha$-escort sampling field while paying a quadratic dissipation cost in $\chi^{2}$ ratio geometry. This single variational principle admits three equivalent interpretations—mirror descent, Hilbert-space projection, and linear response—confirming that OPO is the canonical quadratic proximal response in ratio geometry. The resulting framework has constant Hessian, linear non-saturating gradients, global contraction, and reveals advantage $z$-scoring as a conservation-law projection. Experiments on mathematical reasoning tasks validate that OPO outperforms KL-based methods where gradient saturation limits further improvement.
@@ A.1 Log-Ratio Approximation Error
Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ be the log-ratio, and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$ be the exact ratio deviation. Let $\delta=\|\Delta_{\theta}\|_{\infty}$. For sufficiently small $\delta$ (in particular $\delta<1$, which holds under the small trust-region regime maintained by on-policy anchoring and the $\chi^{2}$ penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$):